Hi, I would appreciate help with a problem from my Algorithms class. Thanks!

*1- This is the Euclid algorithm that computes the greatest

common divisor of two non-negative integers a and b assuming

that a > b >= 0.

Euclid(a, b)       

if b = 0                         

then return a                 

else Euclid(b, a mod b)

Is this algorithm efficient? If so, what is the reason? If not, why not?

Explain your opinion.

*2- Prove that both versions of the knapsack problem possess the optimal

substructure property.

*3- Prove that 3 CNF Satisfiability problem is reducible to the Halting Problem.
Answer #1

1. Yes this algorithm is efficient consider the below methods.:-

The Euclidean Algorithm for finding GCD(A,B) is as follows:

  • If A = 0 then GCD(A,B)=B, since the GCD(0,B)=B, and we can stop.
  • If B = 0 then GCD(A,B)=A, since the GCD(A,0)=A, and we can stop.
  • Write A in quotient remainder form (A = B⋅Q + R)
  • Find GCD(B,R) using the Euclidean Algorithm since GCD(A,B) = GCD(B,R)

Example:

Find the GCD of 270 and 192

  • A=270, B=192
  • A ≠0
  • B ≠0
  • Use long division to find that 270/192 = 1 with a remainder of 78. We can write this as: 270 = 192 * 1 +78
  • Find GCD(192,78), since GCD(270,192)=GCD(192,78)

A=192, B=78

  • A ≠0
  • B ≠0
  • Use long division to find that 192/78 = 2 with a remainder of 36. We can write this as:
  • 192 = 78 * 2 + 36
  • Find GCD(78,36), since GCD(192,78)=GCD(78,36)

A=78, B=36

  • A ≠0
  • B ≠0
  • Use long division to find that 78/36 = 2 with a remainder of 6. We can write this as:
  • 78 = 36 * 2 + 6
  • Find GCD(36,6), since GCD(78,36)=GCD(36,6)

A=36, B=6

  • A ≠0
  • B ≠0
  • Use long division to find that 36/6 = 6 with a remainder of 0. We can write this as:
  • 36 = 6 * 6 + 0
  • Find GCD(6,0), since GCD(36,6)=GCD(6,0)

A=6, B=0

  • A ≠0
  • B =0, GCD(6,0)=6

So we have shown:

GCD(270,192) = GCD(192,78) = GCD(78,36) = GCD(36,6) = GCD(6,0) = 6

GCD(270,192) = 6